OpenStudy (anonymous):

On a buying trip in Los Angeles, Rosaria Perez ordered 120 pieces of jewelry: a number of bracelets at $4 each and a number of necklaces at $13 each. She wrote a check for $1110 to pay for the order. How many bracelets and how many necklaces did Rosaria purchase?

OpenStudy (darthsid):

This is a typical word problem for a system of equations in two variables. Let the number of bracelets be 'b' Let the number of necklaces be 'n' The first line says she ordered 120 pieces, which means there are 120 bracelets and necklaces. This can be written as the equation: b + n = 120 The second line says 'a number of bracelets at $4 each' So, the total value for 'b' number of bracelets is 4*b Similarly, according to the line "necklaces at $13 each" the total value of 'n' necklaces is 13*n Now, the total value is 4b + 13n Which is the same as what she paid, $1110 So, you get two equations: b+n = 120 4b+13n=1110 Solve these two to get your answer!

OpenStudy (anonymous):

55 bracelets and 65 necklaces?

OpenStudy (darthsid):

that's not the answer I am getting, can you post your working?